the north wall of a barn is a rectangle 50 feet high and 50 feet long. There are five windows in the wall, each 5 by 7 feet. In a gallon of paint will cover 400 square feet, how many gallons of paint must the painter buy to paint the wall.
Answer by Cromlix(4381)

Total area of wall = 50 * 50 = 2500 ft^2
Total area of 5 windows 5 * 7 ft
= 175 ft^2
Wall area - window area = 2325 ft^2
1 gallon of paint will cover 400 ft^2
2325/400 = 5.8 gallons
Painter should buy 6 gallons to paint
the north wall.
Hope this helps.
:-)